OpenStudy (anonymous):

0.8x-0.1(16-5x)=-0.2(x-7) simplify the answer

ganeshie8 (ganeshie8):

multiply 10 thru out

ganeshie8 (ganeshie8):

that fixes the decimals, and makes the equation look less messy

ganeshie8 (ganeshie8):

0.8x-0.1(16-5x)=-0.2(x-7) multiply both sides with 10 10[0.8x-0.1(16-5x)] = 10[-0.2(x-7)] 8x - 1(16-5x) = -2(x-7) 8x - (16-5x) = -2(x-7)

ganeshie8 (ganeshie8):

see if u can do the rest.. .

OpenStudy (anonymous):

no